OpenStudy (bibiannaxx):

To get on the top players’ list, Don needs to have a minimum average score of 225 after playing four games. His scores on his first three games were 192, 214, and 250. What is the minimum score Don needs to earn on his fourth game?

OpenStudy (mathmale):

Keep in mind that arithmetic averages are found by summing up the numbers and dividing the sum by the count of these numbers. The average of 2, 3, 5, 7 and 9 is (2+3+5+7+9)/5.

OpenStudy (awesomedude2078):

(192 + 214 + 250 + x)/4 = 225

OpenStudy (mathmale):

Let m = the minimum score that this guy has to achieve on the 4th game. Then the average formula works out to \[\frac{ 192+214+250+m }{ 4}\]

OpenStudy (mathmale):

and since this is a minimum, you need to write in \[\ge 225.\]

OpenStudy (mathmale):

Solve this inequality for m.

OpenStudy (mathmale):

awesome: NO direct answers.

OpenStudy (awesomedude2078):

yep

OpenStudy (mathmale):

NO direct answers, please.

OpenStudy (bibiannaxx):

thank you guys :)

OpenStudy (awesomedude2078):

you're welcome